Find the future value of the following

ordinary

simple

annuity.

Periodic Payment

Payment Interval

Term

Interest Rate

Conversion Period

$444.00444.00

3 months3 months

3.753.75

years

77 %

quarterlyquarterly
Answer #1

Future Value=Periodic Payment/(rate/4)*((1+Rate/4)^(4*t)-1)=444/(7%/4)*((1+7%/4)^(4*3.75)-1)=7541.09551